1/18/00 Owl Shift


Subject: 1/18/00 Owl Shift
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:29:51 CST


 
              Shift Summary for OWL shift, 1/18/2000 (Tue.)

On Shift: Peter Shanahan, Carlos Escobar. Rick Kessler and Wah, plus several
others.

 KAMI TEST IN PROGRESS.
 
Hours of beam available at all intensities: 5.5 @ 2E12

Accounting of lost time & intensity:

      Hall open for change-over for KAMI test running.

========================================================================
Runs ended this shift:
     
       
       15568 KTeV Pedestal Run.
       15570 KAMI_MUON: 1st Kami muon run! No taoe, no disk.
       15571 KTeV Pedestal Run. No tape, write to disk only.
       15572 KAMI_MUON. No tape, no disk.
       15573 KAMI_MUON. No tape, no disk.
       15578 KAMI_MUON. Written to tape and disk for UCAL.

Tapes brought to FCC this shift:
   
       None.

========================================================================
Accesses during shift:

       One access after finishing open access: PS, YW and CE went down to
pull out NM4AN crash button and open NM3 gate valve.

========================================================================
Overview of shift:

       KAMI test running: MUON RUN

========================================================================
Description of Shift:

 00:30 RSK Tests UCAL channel map by raising gain x10 on 1
                   chan.
 00:30 VJ moves TRDs out.
 01:30 15570 MUON_KAMI! No tape, no disk - just for signal timing.
 01:45 Load tapes QKE201-209, QKB169. Tapes QKE182-190 are
                   unaccounted for. Tapes QKE191-200 are deliberately
                   skipped.
 02:20 RSK: HG-right (=UCAL 16) is 20ns too late for L1.
        15571
        15572 More muon runs for tuning UCAL HV.
 04:15 15578 Muon run to disk and tape. DCALIGN didn't run properly.
 05:05 Access to open gate valve and pull crash buttons
        15581 DCALIGN finally sorted out (namelist problems). HN
                   okays update. There were major DC4 changes due to
                   TRDs being moved to out position, as expected.
 05:30 15582 Bit problem in UCAL ADC? Spikey ADC distribution, but
                   swapping doesn't help.
 06:48 15585 KAMI_LAMPPI - 1st hadron run. Rates seem too low...

========================================================================
Follow up Notes on Previous Shift

   We paged HN and installed new dcalign constants.

========================================================================
Items to Follow up on Next Shifts
 
Take Kami data (hadrons).



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:29:53 CST